Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
Gherardini - 8:50 Friday 16 October 2020 (49669) Print this report
Operator Report - initial status
State of Virgo: Unlocked.

ITF Mode: Upgrading.

Quick Summary: IMC locked; SWEB & SNEB in-air; all suspensions loops closed with the exception of SR local controls; SR-BS Valve close; WE_UPS1 not working correctly; TCS CO2 lasers off; INJ_MAIN in pause, all the other automation nodes OFF.

Activities ongoing since this morning: no activity known so far in the control room.
Images attached to this report
Comments to this report:
ruggi - 9:18 Friday 16 October 2020 (49670) Print this report

The red flags warning that SR local controls and guardian are open do not tell the truth: SR is in standard state, and it was like that most of the time that it has been reported the contrary. Investigation in progress.

Images attached to this comment
Boschi - 17:32 Friday 16 October 2020 (49678) Print this report

The SR red flags-problem was due to the recent update of the DSP code in the boards connected to mirror and marionette. The list of gnames has to be refreshed using the relative gain client ('Update list' function in the menu). The problem is now fixed.

Search Help
×

Warning

×